# Differentially Private Sampling via Reveal-or-Obscure

### v2

@@ III Reveal-or-Obscure (ROO)
$\left|\hat{P}_{x^{n}}(x)-\hat{P}_{\tilde{x}^{n}}(x)\right|\leq\frac{1}{n}.$ (13)
@@
$\frac{\Pr\{Y=y|x^{n}\}}{\Pr\{Y=y|\tilde{x}^{n}\}}\leq e^{\epsilon}.$ (14)
@@
$\frac{\Pr\{Y=y|x^{n}\}}{\Pr\{Y=y|\tilde{x}^{n}\}}$ $=\frac{\frac{q}{k}+(1-q)\hat{P}_{x^{n}}(y)}{\frac{q}{k}+(1-q)\hat{P}_{\tilde{x}^{n}}(y)}$ (15)
$\leq\sup_{p\in[0,1]}\frac{\frac{q}{k}+(1-q)\left(p+\frac{1}{n}\right)}{\frac{q}{k}+(1-q)p}$ (16)
$=\sup_{p\in[0,1]}1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}+(1-q)p}$ (17)
$\leq 1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}},$ (18)
where, in (15), we substitute (11); in (16), we use the property of the empirical probabilities stated in (13), denoting $\hat{P}_{\tilde{x}^{n}}(y)$ with $p$ for notational simplicity; in (18), the supremum is obtained when $p=0$.
Therefore, ROO satisfies $\epsilon$-DP guarantee if the right side of (18) is bounded by $e^{\epsilon}$, i.e.,
@@
$\quad\quad+(1-q)\sum_{x^{n}\in\mathcal{X}^{n}}\hat{P}_{x^{n}}(y)\Pr\left\{X^{n}=x^{n}\right\}$ (21)
@@
where we rearrange and substitute (23) to obtain the expression in (25). Note that, for a convex objective function, the maximum is achieved at its corner points. Hence, the distribution $P$ that maximizes the $d_{TV}$ between the uniform distribution on $[k]$ and the input distribution $P$ must be one of the corner points of the $k$-dimensional probability simplex $\mathcal{P}$, e.g. $P=\{1,0,\dots,0\}$. The maximum $d_{TV}$ is then computed as
$\max_{P(y)}d_{TV}\left(\frac{1}{k},P(y)\right)$ $=\max_{P(y)}\frac{1}{2}\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-P(y)\right|=1-\frac{1}{k}.$ (28)
@@
Substituting the above into (19) and rearranging, we obtain sample complexity in (9). ∎
The expression in (9) reveals that Algorithm 1 requires substantially fewer samples than previous mechanisms. The following lemma formalizes this comparison.
Lemma 1.
The proof of Lemma 1 is given in Appendix A. In fact, the sampling complexity of Algorithm 1 is exponentially better in terms of $\epsilon$ compared to that of Raskhodnikova et al. [8] that we summarize in (5), and Cheu and Nayak [9] summarized in (6). This would appear to violate the lower bound of [8] on the sampling complexity; the reason it does not is that their lower bound only applies for $\epsilon<1$.
IV Data-Specific Reveal-or-Obscure (DS-ROO)
For Algorithm 1 in Section III, we fix $q$ in order to achieve $\epsilon$-DP for any possible dataset. From (18) in the proof of Theorem 1, we observe that the supremum of ratio of probabilities is achieved by setting $p=0$, where $p$ represents the empirical probability of one of the elements of the alphabet. This corresponds to the case where one of two neighboring datasets under consideration is entirely missing an element of the alphabet, but this element is present in the other dataset. In this section, we show that for datasets where each element of the alphabet appears reasonably often, it is possible to sample more efficiently than Algorithm 1. By making the obscuring probability $q$ a function of the smallest empirical probability of the dataset, the accuracy of the sampler can be improved for the same privacy. Let $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor\right\}$ denote the smallest number of times an element of $[k]$ appears in dataset $x^{n}$. That is, $m=n\cdot\min_{x\in[k]}\hat{P}_{x^{n}}(x)$. We note that since $X^{n}$ is a random vector, more generally, one can write a random variable $M$ that is the same function applied to $X^{n}$, for which $m$ is an instantiation.
The modified private sampler—which we call the data-specific reveal-or-obscure (DS-ROO)—is presented in Algorithm 2 and requires a one-time computation of $q_{m}$ which is outlined in steps 1–14. In DS-ROO, the data-specific obscuring probability $q_{m}$ is determined from the value of $m$ associated with the given dataset $x^{n}$. Algorithm 2 first computes the obscuring probability value $q_{m}$ via a recursive procedure; at first sight, it may not be obvious why this procedure is the way it is—its correctness will be demonstrated in the proof that follows. In steps 15–18, Algorithm 2 computes the minimum empirical probability $m$ from the input dataset $x^{n}$, then samples using the pre-computed $q_{m}$ as the obscuring probability. [^1] In fact, the computation of $(q_{m}:m\in\{0,\ldots,\lfloor\frac{n}{k}\rfloor\})$ can be done offline and only once given $k,n,\epsilon$.
Algorithm 2 Data-specific ROO (DS-ROO)
2:Sample $y$
3:Set $u^{\prime}\leftarrow-1+\frac{1}{k}-\frac{1}{n}$
4:Set $v^{\prime}\leftarrow e^{\epsilon}\!\left(\frac{1}{k}-1\right)$
5:Set $w^{\prime}\leftarrow e^{\epsilon}-1-\frac{1}{n}$
6:Compute $q_{0}\leftarrow\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}$
7:for $j=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor$ do
8:  $u_{j}\leftarrow-\frac{j}{n}+\frac{1}{k}-\frac{1}{n}$
9:  $v_{j}\leftarrow e^{\epsilon}\!\left(\frac{1}{k}-\frac{j}{n}\right)$
10:  $w_{j}\leftarrow-\frac{1}{n}-\frac{j}{n}+\frac{j}{n}e^{\epsilon}$
11:  if $j<\frac{n}{k}$ then
12:   $q_{j}\leftarrow\max\!\left\{\,0,\;\frac{u_{j}}{v_{j}}q_{j-1}-\frac{w_{j}}{v_{j}},\;\frac{v^{\prime}}{u^{\prime}}q_{j-1}+\frac{w^{\prime}}{u^{\prime}}\,\right\}$
13:  else if $j=\frac{n}{k}$ then
14:   $q_{j}\leftarrow\max\!\left\{\,0,\;\frac{v^{\prime}}{u^{\prime}}q_{j-1}+\frac{w^{\prime}}{u^{\prime}}\,\right\}$
15:  end if
16:end for
17:Compute $m\leftarrow n\cdot\displaystyle\min_{x}\hat{P}_{x^{n}}(x)$
18:With probability $q_{m}$, choose $y\sim\mathrm{Unif}[1\!:\!k]$
19:Otherwise, sample $i\sim\mathrm{Unif}[1\!:\!n]$ and set $y\leftarrow x_{i}$
20:return $y$
Notably, when $m=0$, the corresponding $q_{0}$ is equivalent to that of Algorithm 1. That is, the case of a dataset that is missing an element of the alphabet ($m=0$) is the worst-case scenario, and so in this case DS-ROO behaves identically to ROO. However, we will show both analytically and empirically that in other cases, DS-ROO can do much better than vanilla ROO.
Fig. 1 shows the plot of $q_{m}$ vs $m$ for a fixed $k=10,n=1000$, and several different values of the privacy parameter $\epsilon$. We observe that $q_{m}$ is non-increasing in $m$. As $m$ increases, the empirical distribution of the dataset gets closer to the uniform distribution, and $q_{m}$ approaches zero. Thus, DS-ROO is less likely to obscure the empirical distribution for larger $m$. In high privacy regimes, $q_{m}$ tends to decrease much slower. On the other hand, for larger $\epsilon$, i.e., low privacy, $q_{m}$ goes to zero much faster. The privacy guarantee provided by DS-ROO is given by the following theorem.
Theorem 2.
DS-ROO (Algorithm 2) is $\epsilon$-differentially private.
The complete proof of Theorem 2 is given in Appendix B. We provide a proof sketch below.
Proof:
For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, let $m$ and $\tilde{m}$ denote the smallest number of times any element of $[k]$ appears in $x^{n}$ and $\tilde{x}^{n}$, respectively. Since the two datasets differ in exactly one entry, there are three possible relationships between $m$ and $\tilde{m}$: (1) $\tilde{m}=m$, (2) $\tilde{m}=m+1$, or (3) $\tilde{m}=m-1$. Considering each of these cases separately and applying Definition 2, we derive the conditions under which DS-ROO satisfies $\epsilon$-DP as
$(u_{m}-v_{m})q_{m}$ $\geq w_{m},$ $\text{for all }m,$ (31)
$(u^{\prime}-v^{\prime})q_{m}$ $\leq w^{\prime},$ $\text{for all }m,$ (34)
$u^{\prime}q_{m+1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $\text{for }m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (35)
$u^{\prime}q_{m-1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $\text{for }m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor.$ (36)
where $u^{\prime},v^{\prime},w^{\prime},u_{m},v_{m}$, and $w_{m}$ are as defined in Algorithm 2. Additionally, for $q_{m}$ to be a valid probability, we must also have $0\leq q_{m}\leq 1$ for all $m$. We show that the following expression for $q_{m}$ satisfies the above conditions:
$q_{m}=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}},\frac{v^{\prime}}{u^{\prime}}q_{m-1}+\frac{w^{\prime}}{u^{\prime}}\right\}.$ (37)
Here, the first term ensures $q_{m}$ is non-negative. The two remaining terms are obtained by rearranging (33) and (35). The quantity $q_{0}$ corresponds to the obscuring probability of the worst-case setting ($m=0$) and is equal to the $q$ value used in the original version of ROO, given in (10).
We subsequently show that:
  1. 1.
    $q_{m}$ is non-increasing in $m$.
  2. 2.
    The function $q_{m}$ in (37) satisfies the conditions in (31), (32), (34), and (36).
Therefore, Algorithm 2 is $\epsilon$-differentially private. ∎
The accuracy or utility guarantee of Algorithm 2 is given by the following theorem.
Theorem 3.
Assume the distribution $P$ satisfies $P(x)\geq\gamma$ for all $x\in[k]$, where $\gamma>0$. Given $k\geq 2$, $\epsilon>0$, and
$q_{0}=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)},$ (38)
DS-ROO (Algorithm 2) satisfies
$d_{TV}(P,Q)\leq\bigl(q_{m_{0}}+q_{0}\delta\bigr)\!\left(1-\frac{1}{k}\right),$ (39)
where $\delta=ke^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}$, and $q_{m_{0}}$ is the obscuring probability produced by Algorithm 2 at $m_{0}$. Furthermore, if there exists an integer $m_{0}$ such that $n,k,\epsilon,\gamma,m_{0}$ satisfy
$m_{0}$ $\geq\max\left\{1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right),\frac{2}{e^{\epsilon}-1}\right\},$ (40)
$n$ $\geq\max\left\{\frac{m_{0}}{\gamma},\frac{k-1+\sqrt{(k-1)^{2}+4k(e^{\epsilon}-1)^{2}}}{2(e^{\epsilon}-1)^{2}},2k\right\},$ (41)
then DS-ROO (Algorithm 2) is $\alpha$-accurate with
$\alpha=q_{0}(k-1)e^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}.$ (42)
The complete proof of Theorem 3 is given in Appendix C. We provide a proof sketch below.
Proof:
Lemma 8 shows that for any $\gamma^{\prime}<\gamma$,
$\Pr\!\left\{M<n\gamma^{\prime}\right\}\leq\delta=ke^{-2n(\gamma-\gamma^{\prime})^{2}},$ (43)
where $M=n\cdot\min_{x\in[k]}\hat{P}_{X^{n}}(x)$ is a random variable denoting the smallest empirical count in the dataset. Thus, the event $\{M<m_{0}\}$ with $m_{0}=n\gamma^{\prime}$ becomes exponentially unlikely as $n$ grows.
The utility of DS-ROO is determined by the total variation distance between the output distribution $Q$ of Algorithm 2 and the true distribution $P$. We can decompose $Q$ according to whether $M\geq m_{0}$ or $M<m_{0}$. Conditioned on $M\geq m_{0}$, the non-increasing obscuring probability $q_{M}$ is at most $q_{m_{0}}$, which implies
$\sum_{y}|Q(y)-P(y)|\leq 2q_{m_{0}}\left(1-\frac{1}{k}\right).$ (44)
When $M<m_{0}$, the mechanism behaves no worse than the ROO sampler, which implies
$\sum_{y}|Q(y)-P(y)|\leq 2q_{0}\left(1-\frac{1}{k}\right)\delta.$ (45)
Combining the two cases, we obtain the total variation distance
$d_{TV}(P,Q)\leq\bigl(q_{m_{0}}+q_{0}\delta\bigr)\!\left(1-\frac{1}{k}\right).$ (46)
The upper bound in (46) highlights the improvement on the utility of DS-ROO over the vanilla ROO algorithm. For ROO, the total variation distance is at most $q\left(1-\frac{1}{k}\right)$, where the obscuring probability $q$ is fixed and independent of the dataset. In contrast, the upper bound in (46) is smaller since DS-ROO operates with a reduced obscuring probability $q_{m}\leq q_{m_{0}}$, and the contribution of $q_{0}$ is attenuated by $\delta$. Note that, $q_{m_{0}}=0$ is a sufficient condition to eliminate the first term in (46). We show in Lemma 9 that $q_{m_{0}}=0$ whenever $m_{0}$ and $n$ satisfy the conditions in (40) and (41). Then the first term in (46) vanishes, and using $\delta=ke^{-2n(\gamma-\gamma^{\prime})^{2}}$ and $\gamma^{\prime}=\frac{m_{0}}{n}$ we obtain
$d_{TV}(P,Q)\leq\alpha=q_{0}(k-1)e^{-2n\left(\gamma-\frac{m_{0}}{n}\right)^{2}}.$ (47)
Corollary 1.
For any distribution $P$ satisfying $P(x)\geq\gamma$ for all $x\in[k]$, and any $m_{0}$ satisfying (40), as $n\to\infty$, DS-ROO (Algorithm 2) achieves $\epsilon$-DP and $\alpha$-accuracy with
$\alpha$ $=q_{0}(k-1)e^{-2n\gamma^{2}+4\gamma m_{0}+o(1)}.$ (48)
This implies that the total variation distance goes exponentially to zero as $n\to\infty$.
V Numerical Results
In this section, we empirically compare the utility of DS-ROO with vanilla ROO and the state-of-the-art private sampling algorithms under same parameters. Recall from Definition 1 that the utility, i.e., accuracy, is measured in terms of total variation distance, and it is upper bounded by $\alpha$. Hence, smaller values of $\alpha$ corresponds to more accurate sampling. For the samplers of [8] and [9], we use the corresponding upper bounds on $d_{TV}$ as stated in Table I. From Theorem 1, we have the utility of ROO
$d_{TV}^{\text{ROO}}=\frac{1-\frac{1}{k}}{1+\frac{n}{k}(e^{\epsilon}-1)}.$ (49)
To evaluate the utility of DS-ROO for given values of $\gamma$, $n$, $\epsilon$, and $k$, we use the upper bound in (46) on the total variation distance. We compute $q_{m}$ for each $m$, and then evaluate (46) for each $m_{0}\in[1,n\gamma)$, which define as $d_{TV}(m_{0})$. Since DS-ROO reduces to vanilla ROO where all the assumptions made for the data-specific regime do not hold, we compute the best upper bound on the total variation of DS-ROO as:
$d_{TV}^{\text{DS-ROO}}(n)=\min\left\{d_{TV}^{\text{ROO}},\min_{m_{0}\in[1,n\gamma)}d_{TV}(m_{0})\right\}.$ (50)
Figure 2 shows the comparison of utility across a range of parameter settings. The top row shows results for fixed $k$ and $\epsilon$ while varying $\gamma$, and the bottom row shows results for fixed $\epsilon$ and $\gamma$ while varying $k$. Across all settings, ROO and DS-ROO outperform prior methods, with DS-ROO consistently achieving equal or better utility than vanilla ROO. For sufficiently large $\gamma$, the total variation distance for DS-ROO decays exponentially in $n$. We also observe that this exponential improvement is robust to variations in $\epsilon$ and $k$.
Figure 2: Comparison of accuracy in terms of $d_{TV}$ vs dataset size $n$ for various differentially private sampling algorithms. Top row: fixed $k,\epsilon$, varying $\gamma$; bottom row: fixed $\epsilon,\gamma$, varying $k$.
VI Conclusion
In this work, we study differentially private sampling algorithms for discrete distributions on a finite alphabet. For the algorithm we call Reveal-or-Obscure (ROO), we establish a strictly lower sampling complexity bound than was known in the literature. In addition, we propose a novel generalization—Data-Specific Reveal-or-Obscure (DS-ROO)—which adaptively selects the probability of obscuring the empirical distribution of the dataset to improve the privacy-utility trade-off. We prove that DS-ROO preserves the same $\epsilon$-DP guarantee as ROO while achieving better accuracy. Our theoretical bounds are supported by empirical evaluation demonstrating that DS-ROO outperforms existing private samplers when the underlying distribution is not in the worst-case configuration. For future work, we aim to explore the practicality of our proposed methods for more complex distribution classes, as well as for generating multiple samples.
Appendix A Proof of Lemma 1
@@
$\quad=\frac{2k}{\alpha\epsilon}\frac{\epsilon}{2(e^{\epsilon}-1)}\left(1-\frac{1}{k}-\alpha\right)<\frac{2k}{\alpha\epsilon}=n^{\prime}.$ (52)
As $\epsilon$ increases, the term $e^{\epsilon}-1$ in the denominator of (9) grows exponentially, resulting in significantly lower $n$ compared to $n^{\prime}$. Secondly, we compare the sampling complexity of ROO with that of [9]. Subtracting  (6) from (9), we have
@@
$\quad\quad=\frac{(\epsilon-e^{\epsilon}+1)(k-1)(1-\alpha)-\alpha\epsilon}{\alpha\epsilon(e^{\epsilon}-1)}<0,$ (56)
@@
Appendix B Proof of Theorem 2
Since $m$ denotes the smallest number of times an element in the alphabet $[k]$ appears in a dataset $x^{n}$, it is at most $\left\lfloor\frac{n}{k}\right\rfloor$ when the empirical distribution is uniform. For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, there are three possible values of $\tilde{m}$: $(1)$ $\tilde{m}=m$, when the different entry in $\tilde{x}^{n}$ is a different element in $[k]$, hence it does not affect the minimum count, $(2)$ $\tilde{m}=m+1$, and $(3)$ $\tilde{m}=m-1$. In order to understand how the behavior of $q_{m}$ changes with $m$, we first analyze the likelihood ratio of the output distributions. Recall from Definition 2 that, for Algorithm 2 to achieve $\epsilon$-DP, the following condition must hold for all possible $x^{n}\sim\tilde{x}^{n}$ pairs:
@@ Appendix B Proof of Theorem 2
If (60) is satisfied at the two endpoints of $\hat{P}_{x^{n}}$, then the condition holds for all $\hat{P}_{x^{n}}$. Now, substituting $\min_{x}\hat{P}_{x^{n}}(x)=\frac{m}{n}$ into (60), we require
$\frac{m}{n}\left(1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})\right)\leq e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-\left(1-q_{\tilde{m}}\right)\frac{1}{n}.$ (61)
Let
$u^{\prime}$ $=-1+\frac{1}{k}-\frac{1}{n},$ (62)
$v^{\prime}$ $=e^{\epsilon}\!\left(\frac{1}{k}-1\right),$ (63)
$w^{\prime}$ $=e^{\epsilon}-1-\frac{1}{n},$ (64)
$u_{m}$ $=-\frac{m}{n}+\frac{1}{k}-\frac{1}{n},$ (65)
$v_{m}$ $=e^{\epsilon}\!\left(\frac{1}{k}-\frac{m}{n}\right),$ (66)
$w_{m}$ $=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}.$ (67)
Considering the three cases of $\tilde{m}$ and simplifying (61), we have the following conditions:
$(u_{m}-v_{m})q_{m}$ $\leq w_{m},$ $\text{for all }m,$ (68)
$u_{m}q_{m+1}$ $\leq v_{m}q_{m}+w_{m},$ $m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (69)
$u_{m}q_{m-1}$ $\leq v_{m}q_{m}+w_{m},$ $m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor.$ (70)
Similarly, substituting $\max_{x}\hat{P}_{x^{n}}(x)=1$ into (60), we require
$1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})\leq e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-\left(1-q_{\tilde{m}}\right)\frac{1}{n}.$ (71)
Considering the three cases of $\tilde{m}$ and simplifying (71), we have three more conditions:
$(u^{\prime}-v^{\prime})q_{m}$ $\leq w^{\prime},$ $\text{for all }m,$ (72)
$u^{\prime}q_{m+1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $m=0,1,\ldots,\left\lfloor\frac{n}{k}\right\rfloor-1,$ (73)
$u^{\prime}q_{m-1}$ $\leq v^{\prime}q_{m}+w^{\prime},$ $m=1,2,\ldots,\left\lfloor\frac{n}{k}\right\rfloor.$ (74)
Therefore, for Algorithm 2 to achieve $\epsilon$-DP, the function $q_{m}$ must satisfy (68)–(70) and (72)–(74). Now, recall the recursive definition of $q_{m}$ from Algorithm 2:
$q_{m}$ $=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}},\frac{v^{\prime}}{u^{\prime}}q_{m-1}+\frac{w^{\prime}}{u^{\prime}}\right\}$ (75)
$=\max\left\{0,f_{1}(m,q_{m-1}),f_{2}(q_{m-1})\right\}$ (76)
where we denote $f_{1}(m,q_{m-1})=\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}$ and $f_{2}(q_{m-1})=\frac{v^{\prime}}{u^{\prime}}q_{m-1}+\frac{w^{\prime}}{u^{\prime}}$. By inspection, (75) ensures that (70) and (73) are satisfied. Note that, since $q_{m}$ is a probability, we must have $0\leq q_{m}\leq 1$ for all $m$. The first term in (75) ensures that $q_{m}\geq 0$. The initial value $q_{0}$ is defined as in (10) is at most $1$; as we will show, $q_{m}$ is non-increasing in $m$. This ensures that $q_{m}\leq 1$.
The remainder of the proof is structured as follows:
  • Lemma 2 shows that (75) satisfies the condition in (68).
  • Lemma 3 shows that $f_{1}(m,q_{m-1})\leq q_{m-1}$.
  • Lemma 4 shows that $f_{2}(q_{m-1})\leq q_{m-1}$.
  • By Lemmas 3 and 4, we have
    $q_{m}=\max\left\{0,f_{1}(m,q_{m-1}),f_{2}(q_{m-1})\right\}\leq\max\{0,q_{m-1}\}=q_{m-1}$ (77)
    where the last equality follows from the fact that $q_{m-1}\geq 0$. This proves that $q_{m}$ is non-increasing in $m$. This is used in the subsequent lemmas.
  • Lemma 5 shows that (75) satisfies the condition in (69).
  • Lemma 6 shows that (75) satisfies the condition in (72).
  • Lemma 7 shows that (75) satisfies the condition in (74).
B-A Auxiliary Lemmas
Lemma 2.
For $\epsilon>0,k\geq 2,n>k$, $m\in\left\{0,1,2,\dots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor\right\}$, and $u_{m},v_{m}$ and $w_{m}$ as defined in (65)–(67), the function $q_{m}$ satisfies the condition

### v3

@@ V Conclusion
$|\hat{P}_{x^{n}}(x)-\hat{P}_{\tilde{x}^{n}}(x)|\leq\frac{1}{n}.$ (22)
@@
$\frac{P(Y=y|x^{n})}{P(Y=y|\tilde{x}^{n})}\leq e^{\epsilon}.$ (23)
@@
$\frac{P(Y=y|x^{n})}{P(Y=y|\tilde{x}^{n})}$ $=\frac{\frac{q}{k}+(1-q)\hat{P}_{x^{n}}(y)}{\frac{q}{k}+(1-q)\hat{P}_{\tilde{x}^{n}}(y)}$ (24)
$\leq\sup_{p}\frac{\frac{q}{k}+(1-q)\left(p+\frac{1}{n}\right)}{\frac{q}{k}+(1-q)p}$ (25)
$=\sup_{p}1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}+(1-q)p}$ (26)
$\leq 1+\frac{(1-q)\frac{1}{n}}{\frac{q}{k}}.$ (27)
Here, in (24), we substitute (20). In (25), we use the property of the empirical probabilities stated in (22), denoting $\hat{P}_{\tilde{x}^{n}}(y)$ with $p$ for notational simplicity. In (27), the supremum is obtained when $p=0$. Therefore, ROO satisfies $\epsilon$-DP guarantee if the right side of (27) is bounded by $\epsilon$, i.e.,
@@
$\quad+(1-q)\sum_{x^{n}\in\mathcal{X}^{n}}\hat{P}_{x^{n}}(y)\Pr\left\{X^{n}=x^{n}\right\}$ (30)
@@
where we obtain (34) by rearranging and substituting (32). Note that, for a convex objective function, the maximum is achieved at its corner points. Hence, the distribution $P$ that maximizes the $d_{TV}$ between the uniform distribution on $[k]$ and the input distribution $P$ must be one of the corner points of the $k$-dimensional probability simplex $\mathcal{P}$, e.g. $P=\{1,0,\dots,0\}$. The maximum $d_{TV}$ is then computed as
$\max_{P(y)}d_{TV}\left(\frac{1}{k},P(y)\right)$ $=\max_{P(y)}\frac{1}{2}\sum_{y\in\mathcal{X}}\left|\frac{1}{k}-P(y)\right|$ (37)
$=1-\frac{1}{k}.$ (38)
@@
Substituting the above into (28) and rearranging, we obtain sample complexity
$n=\frac{k(1-\alpha)-1}{\alpha(e^{\epsilon}-1)}.$ (41)
-B Proof of Lemma 1
@@
$=\frac{2k}{\alpha\epsilon}\frac{\epsilon}{2(e^{\epsilon}-1)}\left(1-\frac{1}{k}-\alpha\right)$ (43)
$<\frac{2k}{\alpha\epsilon}=n^{\prime}.$ (44)
As $\epsilon$ increases, the term $e^{\epsilon}-1$ in the denominator of (8) grows exponentially, resulting in significantly lower $n$ compared to $n^{\prime}$. Secondly, we compare the sampling complexity of ROO with that of [1]. Theorem 12 of [1] states that for any $\epsilon>0$ and $\alpha\in(0,1)$, the SubRR algorithm is $\epsilon$-DP and performs $\alpha$-sampling for distributions over $[k]$ with sample complexity
$\frac{1}{\alpha\epsilon}(k-1)(1-\alpha)\leq\frac{k}{\alpha\epsilon}.$ (45)
Subtracting the left side of (45) from (8), we have
@@
$\quad\quad=\frac{(\epsilon-e^{\epsilon}+1)(k-1)(1-\alpha)-\alpha\epsilon}{\alpha\epsilon(e^{\epsilon}-1)}$ (49)
$\quad\quad<0,$ (50)
@@
-C Proof of Theorem 2
Since $m$ denotes the smallest number of times an element in the alphabet $[k]$ appears in a dataset $x^{n}$, it is at most $\left\lfloor\frac{n}{k}\right\rfloor$ when the empirical distribution is uniform. For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, there are three possible values of $\tilde{m}$: $(1)$ $\tilde{m}=m$, when the different entry in $\tilde{x}^{n}$ is a different element in $[k]$, hence it does not affect the minimum count, $(2)$ $\tilde{m}=m+1$, and $(3)$ $\tilde{m}=m-1$. In order to understand how the behavior of $q_{m}$ changes with $m$, we first analyze the likelihood ratio of the output distributions. Recall from Definition 2 that, for Algorithm 2 to achieve $\epsilon$-DP, the following condition must hold for all possible $x^{n}\sim\tilde{x}^{n}$ pairs,
@@
Assuming $q_{m}$ and $q_{\tilde{m}}$ do not change significantly for $x^{n}\sim\tilde{x}^{n}$, we require
$1-q_{\tilde{m}}<e^{\epsilon}(1-q_{m})\Rightarrow 1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})<0.$ (55)
Then. the inequality in (54) becomes
$\hat{P}_{x^{n}}(x)\geq\frac{e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-(1-q_{\tilde{m}})\frac{1}{n}}{1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})}.$ (56)
If (56) holds for all $x\in\mathcal{X}$, it must also hold for the minimum value of $\hat{P}_{x^{n}}(x)$, i.e.,
$\frac{m}{n}\geq\frac{e^{\epsilon}\frac{q_{m}}{k}-\frac{q_{\tilde{m}}}{k}-(1-q_{\tilde{m}})\frac{1}{n}}{1-q_{\tilde{m}}-e^{\epsilon}(1-q_{m})}.$ (57)
Simplifying (57), we arrive at the condition
$\left(-\frac{m}{n}+\frac{1}{k}-\frac{1}{n}\right)q_{\tilde{m}}$ $\leq e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right)q_{m}-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon},$ (58)
$\Rightarrow u_{m}q_{\tilde{m}}$ $\leq v_{m}q_{m}+w_{m},$ (59)
where $u_{m},v_{m}$ and $w_{m}$ are as defined in (10)–(12). Now, for $\tilde{m}=m$, we can replace $q_{\tilde{m}}$ with $q_{m}$ to obtain
@@
Since the denominator is always negative, the right side of (60) will be positive only when the numerator is also negative, i.e.,
$w_{m}=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon}<0\Rightarrow m<\frac{1}{e^{\epsilon}-1}.$ (61)
For $\tilde{m}=m+1$ and $\tilde{m}=m-1$, we obtain two more inequalities,
$u_{m}q_{m+1}\leq v_{m}q_{m}+w_{m},$ (62)
$u_{m}q_{m-1}\leq v_{m}q_{m}+w_{m}.$ (63)
Therefore, for Algorithm 2 to achieve $\epsilon$-DP, the function $q_{m}$ must satisfy the inequalities in (60), (62), and (63). Defining the initial value $q_{0}$ as in (9), and assuming (63) to be an equality for $m\geq 1$, we arrive at the expression
$q_{m}=\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}.$ (64)
Note that, since $q_{m}$ is a probability, we must have $0\leq q_{m}\leq 1$ for all $m$. Therefore, (64) is valid only when the right side is positive. The final form of $q_{m}$ is thus
$q_{m}=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}\right\}.$ (65)
The proofs of (64) satisfying the conditions in (60) and (62) are provided in Lemma 3 and 4. The proof that the function $q_{m}$ is non-increasing is provided in Lemma 2. When $m=0$, $q_{m}$ is equal to that of the fixed $q$ of Algorithm 1. As $m$ increases, the input distribution gets closer to $\text{Unif}[1:k]$, and Algorithm 2 obscures the output distribution with smaller probability.
-D Additional Proofs
Lemma 2
For any $k\geq 2,\epsilon>0$, and datasets of size $n$ where $n>k$, the function $q_{m}$ is non-increasing.